ROTATIONALLY INDUCED STABLE QUANTUM MECHANICAL STATES

\underline{J. Konarski}, Faculty of Chemistry, A.Mickiewicz University of Poznań, Grunwaldzka 6, 60-780 Poznań, Poland.

A new form of potential describing the interaction between atoms at a long distances will be presented. Introduction of the interaction of induced dipole moment with an electric charge into an effective potential describing either intermolecular or intercomplex interaction leads to quantum mechanical states whose stability depends on rotational quantum number J. Forms of the potential and its eigenvalues will be discussed during the presentation.